OpenStudy (anonymous):

tawny found $1.95 in quarters and dimes in the couch cushions of her couch there were four times as many dimes as there were quarters how many dimes and quaters did she find

OpenStudy (anonymous):

if quarters are x, then dimes are 4x. you know \[10\times 4x+25x=195\] solve for x

OpenStudy (anonymous):

hmm this is not giving me a whole number answer...

OpenStudy (anonymous):

answer is 3 quarters and 12 dimes because you get 1.20 +.75=1.95

OpenStudy (anonymous):

oh yes it does give me a whole number asnswer

OpenStudy (anonymous):

\[40x+25x=195\] \[65x=195\] \[x=\frac{195}{65}=3\]

OpenStudy (anonymous):

3 quarters and therefore 12 dimes

OpenStudy (anonymous):

if number of quarters is x then number of dimes is 4x equation should be 0.25x + 0.40x = 1.95 => 0.65x = 1.95 => x = 1.95/0.65 = 3

OpenStudy (anonymous):

so 3 quarters and 12 dimes